28 January 2008

RHIC Running:  Deuteron-Gold running ended this morning at 8:00AM after 12 hours of APEX.  Switch-over to  polarized proton running is underway.  Beam is expected in RHIC tonight. If it is not possible to complete the power supply work today, there will be limited work in the AGS to improve polarization through the night, and finish the RHIC PS work on Tuesday.  The main tasks include switching the polarity of Q7 and the sextupoles.. 

PHENIX had problems over the weekend with a failure of a BBC temperature interlock that disabled all of their triggers.  A first attempt to reset the crate mid-day Saturday did not cure the problem.  A subsequent access during which time the BBC temperature module was replaced finally brought PHENIX back on-line. 

Preparation for pp running includes bringing protons through the AGS at full energy.  Polarization of >50% has been measured in the AGS.  But no beam has been brought out of the AGS.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020